Consider the reaction Cl_2(g) + 3F_2(g) rightarro
Answer #1

if

dCl2/dt = -0.077

relate with respect to F2

mole ratio is

1:3

then the rate is 3x faster

that is

dF2/dt = 3*dCl2/dt

dF2/dt = 3*-0.077 = -0.231 M/s

for B)

same case

but this is PRODUCT so this must be positive

dClF3/dt = +2*(0.077) = 0.154 M